OpenStudy (anonymous):

HELP PLEASE!!!! MEDAL & FAN!!! Josh marked 12% of the e-mails in his Inbox as Important. If he has 125 e-mails in his Inbox, how many did he mark as Important?

OpenStudy (anonymous):

a.) 0.12n = 125 b.) 125n = 0.12 c.) n = 0.12(125) d.) \[n= 0.12\div 125\]

OpenStudy (shelld93):

the formula is \[\frac{ total }{ 100 } \] times the percentage

OpenStudy (shelld93):

\[\frac{ 125 }{ 100 } \times 12\]

OpenStudy (shelld93):

= 15

OpenStudy (shelld93):

He marked 15 e-mails as important

OpenStudy (anonymous):

thanks!! ur a lifesaver!! can u help with more??

OpenStudy (shelld93):

possibly. depends how in depth they go,

OpenStudy (anonymous):

there the same as that one!! i am just checking my answers, i will tag u in them, if thats ok. we r not allowed to to ask more than one question on one post